# AdaAugment: A Tuning-Free and Adaptive Approach to Enhance Data Augmentation

### v2

@@ Abstract
Data augmentation (DA) is widely employed to improve the generalization performance of deep models. However, most existing DA methods use augmentation operations with random magnitudes throughout training. While this fosters diversity, it can also inevitably introduce uncontrolled variability in augmented data, which may cause misalignment with the evolving training status of the target models. Both theoretical and empirical findings suggest that this misalignment increases the risks of underfitting and overfitting. To address these limitations, we propose AdaAugment, an innovative and tuning-free Adaptive Augmentation method that utilizes reinforcement learning to dynamically adjust augmentation magnitudes for individual training samples based on real-time feedback from the target network. Specifically, AdaAugment features a dual-model architecture consisting of a policy network and a target network, which are jointly optimized to effectively adapt augmentation magnitudes. The policy network optimizes the variability within the augmented data, while the target network utilizes the adaptively augmented samples for training. Extensive experiments across benchmark datasets and deep architectures demonstrate that AdaAugment consistently outperforms other state-of-the-art DA methods in effectiveness while maintaining remarkable efficiency.
Data Augmentation
1 Introduction
@@
Figure 1: Comparison between traditional DA methods and adaptive DA methods.
Data augmentation (DA) is a crucial and highly effective technique to alleviate overfitting in the training of deep neural networks (Xu et al., 2023; Yang et al., 2022), resulting in more generalized and reliable models. However, it is important to note that existing DA methods primarily rely on the utilization of totally random or manually devised augmentation magnitudes during training, which can lead to suboptimal training scenarios and require more significant engineering efforts of human experts to tune the hyper-parameters given a new dataset. For instance, information deletion-based DA methods (DeVries & Taylor, 2017; Chen et al., 2020; Yang et al., 2023) randomly erase some sub-regions in images to create augmented samples, leading to random augmentation strengths. Additionally, automatic DA approaches (Cubuk et al., 2020; Müller & Hutter, 2021; Lin et al., 2023a) often entail significant computational overhead to optimize augmentation strategies for each specific dataset before the target model training. The magnitudes of these augmentations are predetermined without any adaptation throughout the training process. Consequently, these methods may fail to capture the unique characteristics of each dataset and the evolving state of network training.
@@
In fact, using augmentation magnitudes introduces inherent randomness, which can increase data diversity but also introduces uncontrolled variability into the augmented data. Such randomness in data variability may not optimally align with the evolving training status of deep models, thereby introducing severe side effects to the training process. For instance, during the initial stage of training, when models generally exhibit weaker generalization capabilities, substantial data variability can induce noise and distributional shifts, potentially leading to underfitting phenomena (Soviany et al., 2022). In contrast, in the later stages of training, limited data variability may elevate the risk of overfitting (Yang et al., 2023). Consequently, this misalignment in the current DA methodologies increases the risks of both underfitting and overfitting, which can ultimately have adverse effects on the generalization performance of models (Gong et al., 2021; Yang et al., 2024). This is a challenge often overlooked in the field.
@@
To tackle these challenges, it is crucial to adopt an adaptive modulation of augmented data variability based on real-time feedback from the target models. As illustrated in Fig. 1, we present a comparison between traditional DA and adaptive DA mechanisms. In contrast to conventional approaches that rely on random or predefined augmentation magnitudes, adaptive DA methods dynamically adjust the magnitudes of DA operations based on real-time feedback from the target network. Notably, adaptive DA focuses on optimizing augmentation strengths rather than the specifics of the operations themselves. This adaptive strategy offers an effective countermeasure to mitigate the aforementioned challenges, thereby unleashing the full potential of DA and enhancing the generalization capabilities of deep models. One straightforward solution to achieve adaptive DA is to create a formal measure that indicates the real-time learning status of each training sample. This measure can then be used to determine the appropriate level of augmentation strengths. Nevertheless, theoretical analyses have identified that determining the overfitting or underfitting risks of a learning algorithm is undecidable (Bashir et al., 2020; Sehra et al., 2021; Li et al., 2024). Consequently, formulating a definitive measure of learning status for training samples remains a formidable challenge.
@@
To address the limitations and challenges above, in this work, we propose AdaAugment, an innovative and tuning-free Adaptive data Augmentation method. AdaAugment leverages a reinforcement learning algorithm to adaptively determine specific augmentation magnitudes for each training sample without relying on any hand-crafted measures. Central to AdaAugment is a dual-model architecture: the policy network and the target network. The policy network learns the policy that determines the magnitudes of augmentation operations for each training sample based on the real-time feedback from the target network during training. The target network simultaneously utilizes these adaptively augmented samples for training. Both networks are jointly optimized, eliminating the need for separate re-training of the target network and enhancing the practical feasibility of our approach. The learned policy adjusts the variability within the augmented data to align with the learning status of the target model, thereby optimizing the introduced variability into the augmented data. Specifically, in the training of the policy network, we estimate the risks of underfitting and overfitting by deriving losses from fully augmented and non-augmented data, respectively. These two losses are then compared with those derived from data adaptively augmented by AdaAugment, serving as reward signals. Extensive experimental results across various benchmark datasets demonstrate the superior performance of AdaAugment compared to state-of-the-art (SOTA) DA methods. Furthermore, complexity analyses of AdaAugment verify that it merely imposes minimal parameters and computational overhead, highlighting its high efficiency. Therefore, AdaAugment achieves a commendable balance between effectiveness and efficiency, delivering high performance without introducing excessive computational complexity. For example, AdaAugment can achieve more than 1% improvements over other SOTA DA methods when training on Tiny-ImageNet without any architectural modifications or additional regularization. In summary, we highlight our contributions as follows:
@@
    AdaAugment features a dual model architecture that jointly optimizes a policy network and a target network. This facilitates concurrent policy formulation and enhances task performance effectively.
@@
    To the best of our knowledge, this is the first work that identifies that the adaptive DA mechanism can dynamically improve models’ generalization.
@@
    Extensive experiments conducted on multiple benchmark datasets and deep networks show that AdaAugment outperforms existing SOTA DA methods. Importantly, AdaAugment achieves this superior performance while incurring minimal additional training overhead, thus ensuring competitive efficiency.
2 Related Work
2.1 Data Augmentation
@@
Data augmentation has played a crucial role in enhancing the generalization capabilities of deep neural networks. Typically, these methods have primarily relied on augmentations characterized by random or predefined magnitudes to introduce diversity into the training data.
@@
Among these methods, Cutout (DeVries & Taylor, 2017) is one of the most widely used techniques, which randomly masks out one or more square regions within training images. Random Erasing (Zhong et al., 2020) randomly selects a rectangular region within an image and erases its pixels with random values. Similarly, Hide-and-Seek (HaS) randomly hides patches in a training image, improving object localization accuracy and the generalization ability of deep models. GridMask (Chen et al., 2020) employs structured dropping regions within the input images. Since these methods may easily introduce noise and ambiguity into the augmented data, AdvMask (Yang et al., 2023) is proposed to identify classification-discriminative information in images and structurally drop some sub-regions that contain critical points for augmentation. Meanwhile, Mixup (Zhang et al., 2018) and CutMix (Yun et al., 2019) mix random information from two or more images to synthesize the augmented data. However, these DA approaches primarily focus on data transformations and often overlook the training status of models. This oversight makes it challenging to alleviate actual overfitting and underfitting risks accordingly through online adjustments to augmentation strengths.
Figure 2: The general framework of the proposed AdaAugment.
@@
Automatic DA methods, such as AutoAugment (Cubuk et al., 2019), Fast-AutoAugment (Lim et al., 2019), and RandAugment (Cubuk et al., 2020), leverage reinforcement learning or grid search to search existing policies for the optimal combination of DA operations on different image datasets in an offline manner. Similarly, Adversarial AutoAugment (Zhang et al., 2019) utilizes fixed augmentation space and rewards the policy yielding the lowest accuracy, causing the policy distribution to shift towards progressively stronger augmentations throughout training. Meanwhile, TrivialAugment (Müller & Hutter, 2021) employs the same augmentation space obtained by these automatic DA methods and applies only a single augmentation operation to each image during training. MetaAugment (Rajendran et al., 2020) utilizes a static augmentation space, and its augmentation policy network outputs weights for augmented data loss. However, the magnitudes of augmentation policies employed in these methods remain fixed or are randomly sampled during online training, leading to the uncontrollable extent of data transformations. SelectAugment (Lin et al., 2023a) employs hierarchical reinforcement learning to acquire online policies that determine the portion of the augmented data and whether each individual sample should be augmented. However, since the specific augmentation transformation applied to each sample is AutoAugment, Mixup, or CutMix, the DA strengths employed are still uncontrollable. KeepAugment (Gong et al., 2021) is proposed to detect and preserve the salient regions of the images during augmentation.
@@
The work (Fawzi et al., 2016) proposes an adaptive DA method that focuses on identifying small transformations that yield maximal classification loss on the augmented data. The work (Lee et al., 2020) leverages impact modeling via the influence function to learn differentiable DA transformation. The work (Xu & Zhao, 2022) randomly determines the types and magnitudes of DA operations for batch-level data and updates DA’s parameters along the gradient direction of the target network loss. The work (Zhang et al., 2023) proposes AdDA for contrastive learning, which allows the network to adjust the augmentation compositions and achieve more generalizable representations. Adaaug (Cheung & Yeung, 2021) learns class-dependent and potentially instance-dependent augmentations via a differentiable workflow. TeachAugment (Suzuki, 2022) transforms data so that they are adversarial to the target model. However, previous adaptive DA methods typically sample augmentation policies with fixed magnitudes and mainly aim to alleviate overfitting risks. More recently, the work (Yang et al., 2024) evaluates DA’s efficacy using similarity and diversity metrics, revealing their varying importance across datasets, suggesting the merit of adaptively adjusting augmentation strengths for optimal balance. Therefore, in contrast to existing DA methods, AdaAugment estimates both underfitting and overfitting risks during training and adaptively adjusts augmentation magnitudes to reduce both risks during online training.
2.2 Reinforcement Learning
@@
Reinforcement learning (RL) is to learn a sequence of actions within an interactive environment by trial and error that maximizes the expected reward (Le et al., 2022; Li, 2023), which obtains a wide range of applications, such as autonomous driving (Kiran et al., 2022) and recommender systems (Lin et al., 2023b), etc. In the realm of RL algorithms, there are two fundamental categories: value optimization and policy optimization methods. Value optimization methods primarily revolve around the estimation of an optimal value function, which subsequently serves as a basis for deriving the optimal policy (Byeon, 2023). Conversely, policy optimization methods estimate the optimal behavior policy without estimating the value functions (Sutton et al., 1999). Moreover, the actor-critic framework, widely adopted in RL, combines the strengths of both value-based and policy-based RL methods (Shakya et al., 2023). This framework encompasses two components: the actor, which is responsible for learning the policy function, and the critic, which is tasked with evaluating the actor’s chosen actions by estimating the value function. This dual mechanism ensures more stable and efficient learning (Grondman et al., 2012), such as Advantage Actor-Critic (A2C) (Mnih et al., 2016).
3 AdaAugment
@@
The primary objective of AdaAugment is to alleviate underfitting and overfitting risks through an adaptive adjustment of DA strengths during training. This adaptive adjustment can be formalized as a decision-making problem on a per-sample basis, thereby eliminating the need for manually crafted metrics. Fig. 2 illustrates the dual-model framework of AdaAugment: While training the target network, a policy network is simultaneously introduced to dynamically optimize the magnitudes of DA operations in the augmentation process. This dual-model framework jointly optimizes two networks, obviating the necessity for separate re-training of the target network and enabling real-time adjustments of augmentation strengths based on the training progress of the target network. Specifically, the policy network optimizes the magnitudes at the sample level in the current training epoch. Then, in the next epoch, the data augmentation process applies augmentation operations to the training samples with these corresponding magnitudes. The adaptively augmented samples are then utilized for optimizing target network training.
@@
An RL task can be formalized as a Markov Decision Process (MDP), which comprises the following components: a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that represents the probability of transitioning from one state to another by taking actions, a reward function $\mathcal{R}$, a discount factor $\gamma\in[0,1]$, and a time step $T$. Given a state $\boldsymbol{s}\in\mathcal{S}$, an RL agent determines an action $a\in\mathcal{A}$ with the policy $\pi(a\mid\boldsymbol{s})$. With these elements, the objective of an RL task is to find an optimal policy, denoted as $\pi^{*}$, that maximizes the expected cumulative reward within the given MDP framework. Meanwhile, suppose that the training dataset $\mathcal{D}$ consists of $N$ training samples, each of the form $(\boldsymbol{x},\boldsymbol{y})\in\mathcal{D}$. Here, $\boldsymbol{x}$ denotes the original data, and $\boldsymbol{y}$ is a $k$-dimensional vector of 0s and 1s indicating the true label of $\boldsymbol{x}$, where $k$ is the total number of classes. We define the augmentation operation as $e(m,\boldsymbol{x})$, where $e$ is drawn from an augmentation space $\mathcal{E}$, and $m$ corresponds to the magnitude of $e$. Compared to previous works (Cubuk et al., 2020; Müller & Hutter, 2021), our approach adaptively determines the magnitude in the augmentation space $\mathcal{E}$ rather than assigning a predefined value before training.
@@
Since the objective of RL is to dynamically determine the appropriate magnitudes for each sample, the state $\boldsymbol{s}\in\mathcal{S}$ should consider three factors: the inherent difficulty associated with each sample, the current training status of the models (e.g., feature extraction capabilities), and the intensity of augmentation operations w.r.t. the former two factors. This multifaceted consideration of state variables is essential for effective RL. Notably, the feature map plays a vital role in this process by providing feedback from the model (Huang et al., 2023; Zhang et al., 2021), encapsulating both the inherent difficulty of the sample and the model’s real-time feature extraction capability. To illustrate this, as shown in Fig. 2, the state vector $\boldsymbol{s}$ in AdaAugment encodes the feature maps of both the non-augmented sample $\boldsymbol{x}$ and adaptively augmented sample $\tilde{\boldsymbol{x}}$, denoted as $\boldsymbol{s}_{none}$ and $\boldsymbol{s}_{ada}$, respectively. Here, $\boldsymbol{s}_{none}$ captures the inherent difficulty of sample $\boldsymbol{x}$ w.r.t. the current target network training status. Meanwhile, $\boldsymbol{s}_{ada}$ encodes the augmentation operation strength applied to $\tilde{\boldsymbol{x}}$ w.r.t. both the real-time target network status and the inherent sample difficulty. By utilizing this information, AdaAugment effectively aligns data variability with the evolving training status of the target networks.

### v3

@@ Abstract
Data augmentation (DA) is widely employed to improve the generalization performance of deep models. However, most existing DA methods employ augmentation operations with fixed or random magnitudes throughout the training process. While this fosters data diversity, it can also inevitably introduce uncontrolled variability in augmented data, which could potentially cause misalignment with the evolving training status of the target models. Both theoretical and empirical findings suggest that this misalignment increases the risks of both underfitting and overfitting. To address these limitations, we propose AdaAugment, an innovative and tuning-free adaptive augmentation method that leverages reinforcement learning to dynamically and adaptively adjust augmentation magnitudes for individual training samples based on real-time feedback from the target network. Specifically, AdaAugment features a dual-model architecture consisting of a policy network and a target network, which are jointly optimized to adapt augmentation magnitudes in accordance with the model’s training progress effectively. The policy network optimizes the variability within the augmented data, while the target network utilizes the adaptively augmented samples for training. These two networks are jointly optimized and mutually reinforce each other. Extensive experiments across benchmark datasets and deep architectures demonstrate that AdaAugment consistently outperforms other state-of-the-art DA methods in effectiveness while maintaining remarkable efficiency. Code is available at https://github.com/Jackbrocp/AdaAugment.
Index Terms:
Data augmentation, generalization, deep learning, computer vision, automated data augmentation.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Recently, deep neural networks (DNNs) have witnessed remarkable advancements. Despite these advancements, achieving satisfactory performance often necessitates larger network architectures and extensive training data. This is particularly challenging as labeled data are sometimes unavailable and expensive to collect [1, 2, 3]. Thus, the ability of deep learning models to generalize beyond the finite training set remains one of the fundamental problems in deep learning [4, 5]. In this context, data augmentation (DA) emerges as a critical and highly effective technique for enhancing DNN generalization performance [6, 7, 8], leading to the development of more generalized and robust models [9, 10, 11, 12]. Indeed, training large deep models to achieve state-of-the-art (SOTA) performance typically employs strong DA techniques.
However, it is important to note that existing DA methods primarily rely on the utilization of totally random or manually devised augmentation magnitudes during training, potentially leading to suboptimal training scenarios [13]. For instance, DA methods based on information deletion or fusion [14, 15, 16, 17, 18, 19, 20] typically erase or mix random sub-regions in images to generate augmented samples, resulting in random augmentation strengths. Automatic DA approaches [21, 22, 23, 24, 25] often incur significant computational overhead to optimize augmentation strategies for each specific dataset prior to target model training. During training, the magnitudes of the applied augmentations are predetermined without any adaptation throughout the entire training process. Moreover, most adaptive DA methods [26, 27, 28, 25] select augmentation operations with random or predefined magnitudes during online training, primarily aiming to alleviate overfitting risks. Consequently, existing DA methods may fail to capture the unique characteristics of different datasets and the entire evolving state of network training, potentially limiting their effectiveness.
@@ I Introduction
Figure 1: Comparison between traditional DA methods and our adaptive DA mechanism. (a) The typical methods employ augmentation operations with random magnitudes and may not align with the training status of deep models. In contrast, in (b), our AdaAugment adaptively determines augmentation magnitudes for each training sample based on the model’s real-time feedback.
@@
Using random augmentation magnitudes or random operations with fixed magnitudes introduces inherent randomness, which can increase data diversity yet introduce uncontrolled variability into the augmented data. Such randomness in data variability may not optimally align with the evolving training status of deep models, thereby introducing severe side effects to the training process. For instance, during the initial training stages, when models generally exhibit weaker generalization capabilities, substantial data variability can induce noises and distributional shifts, potentially leading to underfitting phenomena [29]. In contrast, in the later training stages, limited data variability may elevate the risk of overfitting [16]. Thus, this misalignment in current DA methodologies increases the risks of both underfitting and overfitting, which can ultimately have adverse effects on the generalization performance of models [30, 9]. This is a challenge often overlooked in the field.
@@
To tackle these challenges, it is crucial to adopt an adaptive modulation of augmented data variability based on real-time feedback from the target models. As illustrated in Figure 1, we present a comparison between traditional DA and adaptive DA mechanisms. Unlike conventional approaches that rely on random or predefined augmentation magnitudes, this adaptive DA mechanism dynamically adjusts the magnitudes of DA operations based on real-time feedback from the target network. Moreover, different from most existing adaptive DA methods, this adaptive DA mechanism emphasizes optimizing augmentation strengths rather than the specifics of the operations themselves. The adaptive strategy significantly reduces the search space and offers an effective countermeasure to directly adjust the variability of the augmented data regardless of the specifics of the augmentation policies. This brings several notable advantages: effectively mitigating the aforementioned challenges, unleashing the full potential of DA, and enhancing the generalization capabilities of deep models. A straightforward approach to achieving adaptive DA is to create a formal measure that reflects the real-time learning status of each training sample. This measure can then be used to determine the appropriate level of augmentation strengths, thereby adapting model training status. Nevertheless, theoretical analyses have demonstrated that determining the model status, e.g., underfitting or overfitting risks of a learning algorithm, is undecidable [31, 32, 33]. Consequently, formulating a definitive measure of learning status for training samples remains a formidable challenge.
@@
To address the aforementioned limitations and challenges, in this work, we propose AdaAugment, a novel and tuning-free Adaptive data Augmentation method for image classification. Instead of relying on any hand-crafted measures, AdaAugment leverages a reinforcement learning algorithm to adaptively determine specific augmentation magnitudes for each training sample. Central to AdaAugment is a dual-model architecture: the policy network and the target network. The policy network learns the policy that determines the magnitudes of augmentation operations for each training sample based on the real-time feedback from the target network during training. The target network simultaneously utilizes these adaptively augmented samples for training. Both networks are jointly optimized, eliminating the need for separate re-training of the target network and enhancing the practical applicability of our approach across diverse datasets. The learned policy adaptively adjusts the variability within augmented data to align with the learning status of the target model, thereby optimizing the introduced variability into the augmented data. Specifically, in the training of the policy network, we estimate the risks of underfitting and overfitting by deriving losses from fully augmented and non-augmented data, respectively. These two losses are then compared with those derived from data adaptively augmented by AdaAugment, serving as reward signals. In this way, AdaAugment is capable of adjusting the emphasis to align with model training status, effectively achieving model- and data-adaptive augmentation without requiring any prior knowledge about specific datasets. Extensive experimental results across various benchmark datasets, including coarse-grained classification datasets CIFAR-10/100 [34], Tiny-ImageNet [35], and ImageNet-1k [36], long-tail classification datasets ImageNet-LT and Places-LT [37], and several fine-grained datasets [38, 39, 40, 41] demonstrate the superior effectiveness of AdaAugment in enhancing model generalization performance. Furthermore, complexity analyses of AdaAugment verify that it merely imposes minimal parameters and computational overhead, highlighting its competitive efficiency. Thus, AdaAugment achieves a commendable balance between effectiveness and efficiency, yielding high performance without introducing excessive computational complexity. For instance, AdaAugment can achieve more than 1% improvements over other state-of-the-art DA methods when training on Tiny-ImageNet without any architectural modifications or additional regularization.
In summary, we highlight our contributions as follows:
@@
    AdaAugment features a dual model architecture that jointly optimizes a policy network and a target network. This facilitates policy formulation and enhances task performance concurrently and effectively.
@@
    AdaAugment employs both model- and data-adaptive DA operations to dynamically adapt model training progress.
@@
    Extensive experiments conducted on multiple benchmark datasets and deep architectures demonstrate that AdaAugment outperforms existing SOTA DA methods with competitive training efficiency. This creates a strong baseline of data augmentation for future research.
II Related Work
II-A Data Augmentation
@@
DA has played a crucial role in enhancing the generalization capabilities of deep neural networks. Existing DA methods could be broadly divided into two categories: methods based on information deletion or fusion [15, 42, 19, 17, 18, 14, 16, 43] and automatic DA methods [21, 22, 23, 24, 26, 25, 27, 28, 44]. Typically, these methods have primarily relied on augmentations characterized by random or predefined magnitudes to introduce diversity into the training data.
@@
Methods based on information deletion or fusion. Among these methods, Cutout [14] is one of the most widely used techniques that randomly masks out one or more square regions within images. Random Erasing [43] randomly selects a rectangular region within an image and erases its pixels with random values. Similarly, Hide-and-Seek (HaS) [45] randomly hides patches in a training image, improving object localization accuracy and the generalization ability of deep models. GridMask [15] employs structured dropping regions within the input images. Since these methods may easily introduce noise and ambiguity into the augmented data, AdvMask [16] identifies classification-discriminative information in images and structurally drops some sub-regions containing critical points for augmentation. Additionally, Mixup [17] and CutMix [18] synthesize augmented data by mixing random information from two or more images. However, these DA approaches primarily focus on data transformations and often overlook the training status of models. This oversight makes it challenging to alleviate overfitting risks accordingly through online adjustments to augmentation strengths.
@@
Automatic DA methods. This group of studies can be broadly divided into fixed augmentation policies and online augmentation policies. Methods based on fixed augmentation policies, such as AutoAugment [21], Fast-AutoAugment [46], Faster-AutoAugment [47], and RandAugment [22] leverage reinforcement learning or grid search to search existing policies for the optimal combination of DA operations on different image datasets in an offline manner. Similarly, Adversarial AutoAugment [48] utilizes a fixed augmentation space and rewards the policies that yield the lowest accuracy, causing the policy distribution to shift towards progressively stronger augmentations throughout training. TrivialAugment [23] employs the same augmentation space obtained by these automatic DA methods and applies a single augmentation operation to each image during training. MetaAugment [49] leverages a static augmentation space, and its augmentation policy network outputs weights for augmented data loss. However, the magnitudes of augmentation policies employed in these methods remain fixed or are randomly sampled during online training, leading to the uncontrollable extent of data transformations [50].
@@
On the other hand, methods based on online policies focus on adaptively selecting augmentation policies during training through various optimization processes [26]. Meanwhile, they primarily aim to alleviate overfitting risks. For instance, the work [26] proposes an adaptive DA method that focuses on identifying small transformations that yield maximal classification loss on the augmented data. SelectAugment [24] employs hierarchical reinforcement learning to acquire online policies that determine the portion of the augmented data and whether each individual sample should be augmented. Since the specific augmentation transformation applied to each sample is AutoAugment, Mixup, or CutMix, the DA strengths employed are still uncontrollable. The work [51] leverages impact modeling via the influence function to learn differentiable DA transformation. The work [52] randomly determines the types and magnitudes of DA operations for batch-level data and updates DA’s parameters along the gradient direction of the target network loss. The work [53] proposes AdDA for contrastive learning, which allows the network to adjust the augmentation compositions and achieve more generalizable representations. KeepAugment [30] detects and preserves the salient regions of the images during augmentation. MADAug [50] trains a model-adaptive policy through a bi-level optimization scheme, minimizing a validation set loss of a model trained using the policy-produced data augmentations. Adaaug [27] learns class-dependent and potentially instance-dependent augmentations via a differentiable workflow. Meanwhile, MADAug and Adaaug search not only the applied magnitudes but also the probability of augmentation and specific augmentation policies. TeachAugment [25] transforms data so that they are adversarial to the target model. More recently, the work [9] evaluates DA’s efficacy using similarity and diversity measures, revealing their varying importance across datasets and suggesting the merit of adaptively adjusting augmentation strengths for an optimal balance. Different from existing DA approaches, our proposed AdaAugment merely optimizes the real-value augmentation magnitudes regardless of specific augmentation policies. This brings several notable advantages: the introduced data diversity can be directly optimized regardless of specific augmentation operations; optimizing real-value magnitudes significantly enhances optimization efficiency; and AdaAugment is capable of adapting DA operations with model training status.
II-B Reinforcement Learning
@@
Reinforcement learning (RL) entails learning a sequence of actions within an interactive environment by trial and error that maximizes the expected reward [54, 55]. RL obtains a wide range of applications, such as autonomous driving [56] and recommender systems [57], etc. In the realm of RL algorithms, there are two fundamental categories: value optimization and policy optimization methods. Value optimization methods primarily focus on the estimation of an optimal value function, which subsequently serves as a basis for deriving the optimal policy [58]. Conversely, policy optimization methods estimate the optimal behavior policy without directly estimating the value functions [59]. Moreover, the actor-critic framework, widely adopted in RL, combines the strengths of both value-based and policy-based RL methods [60, 61, 62]. This framework comprises two components: the actor, responsible for learning the policy function, and the critic, tasked with evaluating the actor’s chosen actions by estimating the value function. This dual mechanism ensures more stable and efficient learning [63], such as Advantage Actor-Critic (A2C) [64].
III AdaAugment
Figure 2: The dual-model framework of the proposed AdaAugment.
@@ III AdaAugment
The primary objective of AdaAugment is to mitigate both underfitting and overfitting risks through adaptively adjusting DA strengths during the training phase. Specifically, it accelerates model convergence in early training stages and alleviates overfitting risks in later training stages. This adaptive adjustment can be formalized as a decision-making problem on a per-sample basis, thereby eliminating the need for manually crafted metrics. Figure 2 illustrates the dual-model framework of AdaAugment: While training the target network, a policy network is simultaneously introduced to dynamically optimize the magnitudes of DA operations within the augmentation process. This dual-model framework facilitates the joint optimization of both networks, eliminating the need for separate re-training of the target network and enabling real-time adjustments of augmentation strengths based on the training progress of the target network. Specifically, the policy network optimizes magnitudes at the sample level in the current training epoch. Subsequently, in the next epoch, the data augmentation process applies augmentation operations to the training samples with these corresponding magnitudes. The adaptively augmented samples are then employed to optimize the target network parameters.
@@
An RL task can be formalized as a Markov Decision Process (MDP), consisting of the following components: a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ representing the probability of transitioning from one state to another by taking actions, a reward function $\mathcal{R}$, a discount factor $\gamma\in[0,1]$, and a time step $T$. Given a state $\boldsymbol{s}\in\mathcal{S}$, an RL agent determines an action $a\in\mathcal{A}$ with the policy $\pi(a\mid\boldsymbol{s})$. With these elements, the objective of an RL task is to find an optimal policy, denoted as $\pi^{*}$, that maximizes the expected cumulative reward within the given MDP framework.
TABLE I: The Augmentation Space $\mathcal{E}$.
Transformation $S_{Max}$ Symmetric
identity - -
auto contrast - -
equalize - -
color 1.9 -
contrast 1.9 -
brightness 1.9 -
sharpness 1.9 -
rotation $30^{\circ}$$\pm$
$\text{translate}_{x}$ 10 $\pm$
$\text{translate}_{y}$ 10 $\pm$
$\text{shear}_{x}$ 0.3 $\pm$
$\text{shear}_{y}$ 0.3 $\pm$
solarize 256 -
posterize 4 -
Suppose that the training dataset $\mathcal{D}$ consists of $N$ training samples, each of the form $(\boldsymbol{x},\boldsymbol{y})\in\mathcal{D}$. Here, $\boldsymbol{x}$ denotes the original data, and $\boldsymbol{y}$ is a $k$-dimensional vector of zeros and ones indicating the true label of $\boldsymbol{x}$, where $k$ is the total number of classes. The augmentation operation is defined as $e(m,\boldsymbol{x})$, where $e$ is randomly drawn from an augmentation space $\mathcal{E}$ and $m$ represents the corresponding magnitude of $e$. Although $e$ is randomly selected from $\mathcal{E}$, the applied magnitudes are adaptively determined by AdaAugment. This effectively adjusts the augmented data variability, achieving a superior balance between the diversity and consistency of the training data, leading to improved model generalization.
We provide the specific details of $\mathcal{E}$ in Table I, which contains widely used image operations and their corresponding maximal magnitudes. Notably, the augmentation space $\mathcal{E}$ closely follows the setting from [23, 21, 22, 46]. Unlike previous works, our approach adaptively determines the magnitude in the augmentation space $\mathcal{E}$ rather than assigning a predefined value before training. In this way, the applied magnitudes are determined by $S_{Max}\times m$, where $S_{Max}$ is the maximum allowable strength of the corresponding transformations. Moreover, for symmetric transformations, the symmetric direction is randomly selected.
@@
Since the objective of RL is to dynamically determine the appropriate magnitudes for each sample, the state $\boldsymbol{s}\in\mathcal{S}$ should consider three factors: the inherent difficulty associated with each sample, the current training status of the models (e.g., feature extraction capabilities), and the intensity of augmentation operations w.r.t. the former two factors. This multifaceted consideration of state variables is essential for effective RL. Notably, the feature map plays a vital role in this process by providing feedback from the model [65, 66], encapsulating both the inherent difficulty of the sample and the model’s real-time feature extraction capability. To illustrate this, as shown in Figure 2, the state vector $\boldsymbol{s}$ in AdaAugment encodes the feature maps of both the non-augmented sample $\boldsymbol{x}$ and adaptively augmented sample $\tilde{\boldsymbol{x}}$, denoted as $\boldsymbol{s}_{none}$ and $\boldsymbol{s}_{ada}$, respectively. The feature maps are obtained from the output of the final linear layer before the classification head in the target network. Here, $\boldsymbol{s}_{none}$ captures the inherent difficulty of sample $\boldsymbol{x}$ w.r.t. the current target network training status. Meanwhile, $\boldsymbol{s}_{ada}$ encodes the augmentation operation strength applied to $\tilde{\boldsymbol{x}}$ w.r.t. both the real-time target network status and the inherent sample difficulty. By utilizing this information, AdaAugment effectively aligns data variability with the evolving training status of the target networks.
TABLE II: Network Architecture of A2C.
index Layer Dimension
Actor 1 Linear (512, 512)
2 Linear (512, 256)
3 Linear (256, 1)
Critic 1 Linear (512, 512)
2 Linear (512, 256)
3 Linear (256, 1)